Determine the concentration of H+ in each solution at 25°C. A solution with pH = 1.0. [HT] = M A solution with pH = 4.0. M
A solution with POH 11.0. M [H]=
Answer #1

PP - 1. 1) H slio pH = -log(H+) 1-0 = -log(H) (ut] - lö = 0.1 [+] 0.1 19 2) pH 4.0 I pH -109 (H+) 4.0 = -109(H+) (ht] = 104m